# A Simple Algorithm for Dynamic Carpooling with Recourse

### v1

@@ 1 Introduction
The problem was first studied by Ajtai et al. [AAN+98] as a special case of the general carpooling problem for hypergraphs posed by Faige and Williams in [FW83] and is motivated by fairness in scheduling. Indeed, the above problem exactly corresponds to a set of drivers who wish to carpool in pairs where each driver wishes to drive roughly half the time. If all the edges of the graph are known beforehand, a simple scheme of directing cycles yields a solution of discrepancy at most one. That said, the more natural setting is arguably when the edges arrive online. In this setting, edges arrive sequentially, and one must irrevocably decide on the orientation of an edge before observing the next edge in the sequence. In this model, [AAN+98] showed that in the presence of an adaptive adversary, any orientation algorithm must have $\Omega(n)$ discrepancy in the worst case.
There are several approaches in the literature for circumventing such a lower bound. One, explored in [AAN+98, GKKS20], considers a stochastic version of the problem, namely when edge arrivals are sampled from a known distribution. In such a setting, [GKKS20] obtain a solution achieving $\mathrm{polylog}(n,m)$-discrepancy. In a recent work, [ALS21] improved this bound by showing that one can achieve discrepancy $O(\log(nm))$ in the presence of any oblivious adversary.

### v2

@@ 1 Introduction
The problem was first studied by Ajtai et al. [AAN+98] as a special case of the general carpooling problem for hypergraphs posed by Fagin and Williams in [FW83] and is motivated by fairness in scheduling. Indeed, the above problem exactly corresponds to a set of drivers who wish to carpool in pairs where each driver wishes to drive roughly half the time. If all the edges of the graph are known beforehand, a simple scheme of directing cycles yields a solution of discrepancy at most one. That said, the more natural setting is arguably when the edges arrive online. In this setting, edges arrive sequentially, and one must irrevocably decide on the orientation of an edge before observing the next edge in the sequence. In this model, [AAN+98] showed that in the presence of an adaptive adversary, any orientation algorithm must have $\Omega(n)$ discrepancy in the worst case.
There are several approaches in the literature for circumventing such a lower bound. One, explored in [AAN+98, GKKS20], considers a stochastic version of the problem, namely when edge arrivals are sampled from a known distribution. In such a setting, [GKKS20] obtain a solution achieving $\mathrm{polylog}(n,m)$-discrepancy. In a recent work, [ALS21, LSS21] improved this bound by showing that one can achieve discrepancy $O(\log(nm))$ in the presence of any oblivious adversary. This was further improved to $O(\sqrt{\log(mn)})$ discrepancy by Kulkarni et al. [KRR24], although their methods do not give an efficient algorithm for computing such a solution.